# Reaction-diffusion models of biological invasion: Open source computational tools, key concepts and analysis

### 07 Mar 2024 v4

@@ 1 Introduction
Figure 1: Biological motivation. (a) Continental-scale invasion of populations of humans [1]. (c) Forest-scale invasion of plants and trees during forest recovery in Puerto Rico [2]. (d) Tissue-scale growth and expansion of population of malignant cancer cells [3, 4]. All three examples have been modelled using simple extensions of the classic Fisher-KPP model. (d) Schematic travelling wave with a smooth front where $u(x,t)>0$ on $0<x<\infty$. (e) Schematic travelling wave with a sharp, well–defined front where $u(x,t)>0$ for $x<L(t)$, with $x(L(t),t)=0$. Images in (a)–(c) used with permission.
@@ 2 Fisher-KPP model: Smooth initial conditions
indicating that the faster travelling the wave speed, the flatter the shape of the wave since we have $\textrm{d}U/\textrm{d}z\to 0^{-}$ as $c\to\infty$ [9].
@@ 3 Porous-Fisher model: Smooth initial conditions
$z(\zeta)=\int U(\zeta)\,\textrm{d}\zeta,$ (33)
@@
Figure 7: Approximate shape of the smooth-fronted travelling wave solution of the PF model given by $c\to\infty$ asymptotic expressions for: (a) $c=3$, and (b) $c=1$. Each approximate solution is compared with the shape of the travelling wave obtained by taking a late-time PDE solution from Figure 5 translates so that $U(0)=1/2$ (solid blue). The upper panels compares the full numerical solution with the $\mathcal{O}(1)$ perturbation approximation $U0(z)$ and the lower panel compares the full numerical solution with the $\mathcal{O}(c^{2})$ perturbation approximation $U_{0}(z)+c^{-2}U_{1}(z)$.
@@ 5 Fisher-Stefan model
Simulations in Figure 9(a)–(b) for $\kappa>0$ lead to constant-speed, constant shape invading travelling wave solutions moving in the positive $x$-direction. The travelling wave speed of these solutions can be estimated numerically like we did in Figure 2 by tracking the position of some density contour and using regression. Since the Fisher-Stefan model is a moving boundary problem, there is a more obvious second approach to determine $c$ as we can use our numerical estimates of $u_{N-1}$ and $u_{N-1}$ to evaluate Equation (48) to give a direct estimate of $\textrm{d}L/\textrm{d}t$ without needing to use front tracking and regression. Using this approach at $t=1,2,3,\ldots,50$ indicates that $\textrm{d}L/\textrm{d}t$ settles to a constant speed that agrees very well with the result obtained by front tracking and regression (not shown). In this Section we simply report estimates of $c$ by evaluating (48) using late-time data. The two invading solutions in Figure 9(a)–(b) indicate that $c$ appears to increase with $\kappa$; we will formalise the relationship between these quantities later when we turn to the phase plane. Interestingly, we have $c\approx 0.36$ for $\kappa=1$ and $c\approx 0.548$ for $\kappa=2$, indicating that these two travelling wave speeds are slower than the minimum travelling wave speed for the Fisher-KPP or Porous-Fisher models. Results in Figure 9(c)–(d) for $\kappa<0$ lead to receding travelling wave solutions with $c<0$. Overall, we see that $c$ is an increasing function of $\kappa$ regardless of whether there is an invading or receding travelling wave; again we note that modelling biological recession is simply not possible using the Fisher-KPP or Porous-Fisher models, whereas in the moving boundary framework it is straightforward to model biological recession.
@@
It is interesting to note that we have made analytic progress for slow travelling wave solutions by developing approximate expressions for certain trajectories in the phase plane rather than working directly to obtain approximate expressions for $U(z)$ like we were able to do for fast travelling wave solutions of the Fisher-KPP and Porous-Fisher models when $c\gg 1$. Our approach here of working in the phase plane is advantageous since we are able to use these perturbation solutions to derive an expression relating $c$ to $\kappa$ through the moving boundary condition. To compare the shape of the travelling wave solution $U(z)$ with our time-dependent PDE solutions, we take our numerical estimates of $U(V)$ and recall that $\textrm{d}U/\textrm{d}z=V$ by definition. Integrating this ODE numerically, again using the trapezoid rule, provides us with estimates of $U(z)$. This approach can be used to determine $U(z)$ either using the numerically generated phase plane trajectory for any value of $c$ or the two-term perturbation approximation for $|c|\ll 1$. Using this approach we computed $U(z)$, and compared the two profiles in the lower panel of each subfigure in Figure 10. Here we see that the perturbation approach leads to very accurate profiles of $U(z)$ for $c=\pm 0.1$. We also note that the late-time PDE solutions in Figure 8 are superimposed with red dashed curves that are indistinguishable from the late-time PDE solution. These red dashed curves are obtained by generating the phase plane for the appropriate value of $c$ and using the $V(U)$ trajectory in the phase plane to estimate the $U(z)$ curve in the same way as described here.
@@
$\dfrac{\textrm{d}^{2}U_{1}(\eta)}{\textrm{d}\eta^{2}}+\dfrac{\textrm{d}U_{1}(\eta)}{\textrm{d}\eta}+U_{0}(\eta)(1-U0(\eta))=0,\quad U_{1}(0)=0,\quad U_{1}(\eta)=0\quad\textrm{as}\quad\eta\to\infty.$ (63)
@@
Before concluding our discussion about the Fisher-Stefan model it is insightful to note that using both the time-dependent PDE model and phase plane tools provides us with opportunities to independently check the accuracy of our numerical algorithms. For example, generating the time-dependent PDE solutions in Figure 9 involves specifying $\kappa$ as an input and observing $c$ as an output. This process involves several different approximations since we must discretise the governing PDE, solve the resulting system of semi-discrete ODEs for a sufficiently long period of time before using the discretised boundary condition to provide estimate $c$. For example, time-dependent solutions in Figure 9(a) involve specifying $\kappa=1$, and out estimate of $c=0.364421881$ is impacted by several choices we have made in implementing our numerical algorithm, such as truncation error. One way of testing the accuracy of our approximation is to repeat the calculation by re-solving the time-dependent PDE model on a finer mesh and examine how our estimate of $c$ depends upon the choice of $\delta\rho$. A different way approach is to use the phase plane where we specify $c$ as an input, generate the trajectory leaving $(1,0)$ along the unstable manifold and to use this trajectory to estimate $V^{\dagger}$, from which we can apply the moving boundary condition, $c=-\kappa V^{\dagger}$ to give $\kappa$ as an output of the phase plane calculations. To demonstrate this approach we generated the phase plane with $c=0.364421881$ as an input, solved for the appropriate trajectory to give $V^{\dagger}=-0.365356934$, and then applied the moving boundary condition to find that $\kappa=0.997440715$, which is within 0.26% of the expected result, $\kappa=1$. This inter-code comparison is convenient since it gives us confidence that both the time-dependent PDE algorithm and the numerical phase plane algorithm are giving consistent results. In addition, if we repeat this inter-code comparison by re-solving the time-dependent PDE model with more mesh points we obtain an improved estimate of $c$ which, when used as an input into the phase plane, leads to an improved estimate of $\kappa$ (results not shown).
@@ 7 Appendix: Porous–Fisher–Stefan model
noting that $\phi_{N}=0$ by definition. This system of $N+1$ coupled nonlinear ODEs can be integrated through time using DifferentialEquations.jl to give estimates of $\phi(\rho_{i},t)$ and $L(t)$. Recalling that $u=+\sqrt{\phi}$, we can plot the solution in terms of $u(x,t)$. Note that the appropriate boundary condition on the moving boundary is to enforce $u(L(t),t)=0$. Computationally, however, it necessary to appropriate this as $u(L(t),t)=\mathcal{U}$, with $\mathcal{U}\ll 1$ otherwise the discretised system of ODEs does not evolve with time. All numerical results in Figure 12 are obtained by setting $\phi(L(t),t)=1\times 10^{-8}$. These time-dependent PDE results are reminiscent of those in Figure 9 for the Fisher-Stefan problem as we see a series of sharp-fronted moving fronts that, in the long-time limit, approach a constant-speed and constant-shape travelling wave. Moreover, we see that setting $\kappa>0$ leads to invading travelling waves with $c>0$, whereas setting $\kappa<0$ we obtain receding travelling wave with $c<0$.
@@
To study the travelling wave solutions of the Porous-Fisher-Stefan model we turn to the phase plane. Working in the travelling wave coordinate $z=x-ct$, we note that travelling wave solutions are translationally invariant in $z$ so we are free to choose $z=0$ to represent the location of the moving boundary which gives
@@
$\psi_{0}(\phi)\dfrac{\textrm{d}\psi_{0}(\psi)}{\textrm{d}\phi}$ $=-2\sqrt{\phi}\left(1-\sqrt{\phi}\right),\quad\psi_{0}(1)=0,$ (81)
@@
$\dfrac{\textrm{d}\psi_{0}(\psi)}{\textrm{d}\phi}$ $=-\dfrac{1}{\sqrt{2\phi}}-\dfrac{2\sqrt{\phi}\left(1-\sqrt{\phi}\right)}{\psi_{0}(\phi)},\quad\psi_{0}(1)=0,$ (88)

### 09 Mar 2024 v5

@@ 1 Introduction
Figure 1: Biological motivation. (a) Continental-scale invasion of populations of humans [1]. (c) Forest-scale invasion of plants and trees during forest recovery in Puerto Rico [2]. (d) Tissue-scale growth and expansion of population of malignant cancer cells [3, 4]. All three examples have been modelled using simple extensions of the classic Fisher-KPP model. (d) Schematic travelling wave with a smooth front where $u(x,t)>0$ on $0<x<\infty$. (e) Schematic travelling wave with a sharp, well–defined front where $u(x,t)>0$ for $x<L(t)$, with $x(L(t),t)=0$.
@@ 2 Fisher-KPP model: Smooth initial conditions
indicating that the faster the travelling wave speed, the flatter the shape of the wave since we have $\textrm{d}U/\textrm{d}z\to 0^{-}$ as $c\to\infty$ [9].
@@ 3 Porous-Fisher model: Smooth initial conditions
$z(\zeta)=\int^{\zeta}U(\zeta^{\prime})\,\textrm{d}\zeta^{\prime},$ (33)
@@
Figure 7: Approximate shape of the smooth-fronted travelling wave solution of the PF model given by $c\to\infty$ asymptotic expressions for: (a) $c=3$, and (b) $c=1$. Each approximate solution is compared with the shape of the travelling wave obtained by taking a late-time PDE solution from Figure 5 translates so that $U(0)=1/2$ (solid blue). The upper panels compares the full numerical solution with the $\mathcal{O}(1)$ perturbation approximation $U_{0}(z)$ and the lower panel compares the full numerical solution with the $\mathcal{O}(c^{2})$ perturbation approximation $U_{0}(z)+c^{-2}U_{1}(z)$.
@@ 5 Fisher-Stefan model
Simulations in Figure 9(a)–(b) for $\kappa>0$ lead to constant-speed, constant shape invading travelling wave solutions moving in the positive $x$-direction. The travelling wave speed of these solutions can be estimated numerically like we did in Figure 2 by tracking the position of some density contour and using regression. Since the Fisher-Stefan model is a moving boundary problem, there is a more obvious second approach to determine $c$ as we can use our numerical estimates of $u_{N-1}$ and $u_{N-2}$ to evaluate Equation (48) to give a direct estimate of $\textrm{d}L/\textrm{d}t$ without needing to use front tracking and regression. Using this approach at $t=1,2,3,\ldots,50$ indicates that $\textrm{d}L/\textrm{d}t$ settles to a constant speed that agrees very well with the result obtained by front tracking and regression (not shown). In this Section we simply report estimates of $c$ by evaluating (48) using late-time data. The two invading solutions in Figure 9(a)–(b) indicate that $c$ appears to increase with $\kappa$; we will formalise the relationship between these quantities later when we turn to the phase plane. Interestingly, we have $c\approx 0.36$ for $\kappa=1$ and $c\approx 0.666$ for $\kappa=3$, indicating that these two travelling wave speeds are slower than the minimum travelling wave speed for the Fisher-KPP or Porous-Fisher models. Results in Figure 9(c)–(d) for $\kappa<0$ lead to receding travelling wave solutions with $c<0$. Overall, we see that $c$ is an increasing function of $\kappa$ regardless of whether there is an invading or receding travelling wave; again we note that modelling biological recession is simply not possible using the Fisher-KPP or Porous-Fisher models, whereas in the moving boundary framework it is straightforward to model biological recession.
@@
It is interesting to note that we have made analytic progress for slow travelling wave solutions by developing approximate expressions for certain trajectories in the phase plane rather than working directly to obtain approximate expressions for $U(z)$ like we were able to do for fast travelling wave solutions of the Fisher-KPP and Porous-Fisher models when $c\gg 1$. Our approach here of working in the phase plane is advantageous since we are able to use these perturbation solutions to derive an expression relating $c$ to $\kappa$ through the moving boundary condition. To compare the shape of the travelling wave solution $U(z)$ with our time-dependent PDE solutions, we take our numerical estimates of $U(V)$ and recall that $\textrm{d}U/\textrm{d}z=V$ by definition. Integrating this ODE numerically, again using the trapezoid rule, provides us with estimates of $U(z)$. This approach can be used to determine $U(z)$ either using the numerically generated phase plane trajectory for any value of $c$ or the two-term perturbation approximation for $|c|\ll 1$. Using this approach we computed $U(z)$, and compared the two profiles in the lower panel of each subfigure in Figure 10. Here we see that the perturbation approach leads to very accurate profiles of $U(z)$ for $c=\pm 0.1$. We also note that the late-time PDE solutions in Figure 8 are superimposed with red dashed curves that are indistinguishable from the late-time PDE solution. These red dashed curves are obtained by generating the phase plane for the approximate value of $c$ and using the $V(U)$ trajectory in the phase plane to estimate the $U(z)$ curve in the same way as described here.
@@
$\dfrac{\textrm{d}^{2}U_{1}(\eta)}{\textrm{d}\eta^{2}}+\dfrac{\textrm{d}U_{1}(\eta)}{\textrm{d}\eta}+U_{0}(\eta)(1-U_{0}(\eta))=0,\quad U_{1}(0)=0,\quad U_{1}(\eta)=0\quad\textrm{as}\quad\eta\to\infty.$ (63)
@@
Before concluding our discussion about the Fisher-Stefan model it is insightful to note that using both the time-dependent PDE model and phase plane tools provides us with opportunities to independently check the accuracy of our numerical algorithms. For example, generating the time-dependent PDE solutions in Figure 9 involves specifying $\kappa$ as an input and observing $c$ as an output. This process involves several different approximations since we must discretise the governing PDE, solve the resulting system of semi-discrete ODEs for a sufficiently long period of time before using the discretised boundary condition to provide estimate $c$. For example, time-dependent solutions in Figure 9(a) involve specifying $\kappa=1$, and our estimate of $c=0.364421881$ is impacted by several choices we have made in implementing our numerical algorithm, such as truncation error. One way of testing the accuracy of our approximation is to repeat the calculation by re-solving the time-dependent PDE model on a finer mesh and examine how our estimate of $c$ depends upon the choice of $\delta\rho$. A different approach is to use the phase plane where we specify $c$ as an input, generate the trajectory leaving $(1,0)$ along the unstable manifold and to use this trajectory to estimate $V^{\dagger}$, from which we can apply the moving boundary condition, $c=-\kappa V^{\dagger}$ to give $\kappa$ as an output of the phase plane calculations. To demonstrate this approach we generated the phase plane with $c=0.364421881$ as an input, solved for the appropriate trajectory to give $V^{\dagger}=-0.365356934$, and then applied the moving boundary condition to find that $\kappa=0.997440708$, which is within 0.26% of the expected result, $\kappa=1$. This inter-code comparison is convenient since it gives us confidence that both the time-dependent PDE algorithm and the numerical phase plane algorithm are giving consistent results. In addition, if we repeat this inter-code comparison by re-solving the time-dependent PDE model with more mesh points we obtain an improved estimate of $c$ which, when used as an input into the phase plane, leads to an improved estimate of $\kappa$ (results not shown).
@@ 7 Appendix: Porous–Fisher–Stefan model
noting that $\phi_{N}=0$ by definition. This system of $N+1$ coupled nonlinear ODEs can be integrated through time using DifferentialEquations.jl to give estimates of $\phi(\rho_{i},t)$ and $L(t)$. Recalling that $u=+\sqrt{\phi}$, we can plot the solution in terms of $u(x,t)$. Note that the appropriate boundary condition on the moving boundary is to enforce $u(L(t),t)=0$. Computationally, however, it necessary to approximate this as $u(L(t),t)=\mathcal{U}$, with $\mathcal{U}\ll 1$ otherwise the discretised system of ODEs does not evolve with time. All numerical results in Figure 12 are obtained by setting $\phi(L(t),t)=1\times 10^{-8}$. These time-dependent PDE results are reminiscent of those in Figure 9 for the Fisher-Stefan problem as we see a series of sharp-fronted moving fronts that, in the long-time limit, approach a constant-speed and constant-shape travelling wave. Moreover, we see that setting $\kappa>0$ leads to invading travelling waves with $c>0$, whereas setting $\kappa<0$ we obtain receding travelling wave with $c<0$.
@@
To study the travelling wave solutions of the Porous-Fisher-Stefan model we turn to the phase plane. Working in the travelling wave coordinate $z=x-ct$, we note that travelling wave solutions are invariant with respect to a translation in $z$ so we are free to choose $z=0$ to represent the location of the moving boundary which gives
@@
$\psi_{0}(\phi)\dfrac{\textrm{d}\psi_{0}(\phi)}{\textrm{d}\phi}$ $=-2\sqrt{\phi}\left(1-\sqrt{\phi}\right),\quad\psi_{0}(1)=0,$ (81)
@@
$\dfrac{\textrm{d}\psi_{0}(\phi)}{\textrm{d}\phi}$ $=-\dfrac{1}{\sqrt{2\phi}}-\dfrac{2\sqrt{\phi}\left(1-\sqrt{\phi}\right)}{\psi_{0}(\phi)},\quad\psi_{0}(1)=0,$ (88)